OpenStudy (sherribelle):

If interest is compounded continuously, the following equation relates the amount of time t it takes for an investment to double in value to the annual interest rate r: 2P=Pert. a. Solve the equation for t in terms of r. b. How long does it take for an investment to double in value compounding continuously if the annual interest rate is 7%? Round your answer to the nearest tenth of a year. c. Solve the equation for r in terms of t. d. What interest rate is necessary for an investment to double in value in 7 years if interest is compounded continuously? Round your answer to the near

OpenStudy (sherribelle):

So far I have: a) t= ln(2)/r + ln(P)/r b) t = 14.3 c) r = ln(2)/t + ln(P)/t

OpenStudy (sherribelle):

Also it's suppose to be 2P=Pe^rt